# JL1-CD: A New Benchmark for Remote Sensing Change Detection and a Robust Multi-Teacher Knowledge Distillation Framework

### v3

@@ Abstract
Change detection (CD) in remote sensing images plays a vital role in Earth observation. However, the scarcity of high-resolution, comprehensive open-source datasets and the difficulty in achieving robust performance across varying change types remain major challenges. To address these issues, we introduce JL1-CD, a large-scale, sub-meter CD dataset consisting of 5,000 image pairs. We further propose a novel Origin-Partition (O-P) strategy and integrate it into a Multi-Teacher Knowledge Distillation (MTKD) framework to enhance CD performance. The O-P strategy partitions the training set by Change Area Ratio (CAR) and trains specialized teacher models on each subset. The MTKD framework then distills complementary knowledge from these teachers into a single student model, enabling improved detection results across diverse CAR scenarios without additional inference cost. Our MTKD approach demonstrated strong performance in the 2024 “Jilin-1” Cup challenge, ranking first in the preliminary and second in the final rounds. Extensive experiments on the JL1-CD and SYSU-CD datasets show that the MTKD framework consistently improves the performance of CD models with various network architectures and parameter sizes, establishing new state-of-the-art results. Code and dataset are available at https://anonymous.4open.science/r/MTKD-A-84B8.
1 Introduction
@@
Figure 1: Timeline of the development of mainstream DL-based CD methods. The numbers “1-5” and “6-10” denote publication periods in 2024 (January-May and June-October, respectively).
@@
  • We evaluate numerous existing algorithms on the JL1-CD dataset, establishing a new benchmark for remote sensing CD by publicly releasing the associated dataset and code. This contribution is anticipated to facilitate further advancements in the field.
@@
2 Related Works
2.1 DL-Based CD
Deep learning has experienced rapid advancements, achieving remarkable success in remote sensing image CD. As illustrated in Figure 1, we present a timeline of the development of mainstream DL-based CD algorithms. Based on the differences in neural network architectures and training paradigms, we briefly review three major categories of DL-based CD methods that are widely adopted in current research.
CNN-Based CD: CNNs serve as the foundation of many early DL-based CD methods and remain widely used today. Daudt et al. [2] proposed three fully convolutional neural network architectures, marking the entry of DL into CD research. This breakthrough spurred numerous subsequent CNN-based solutions. More recently, researchers have also developed efficient CD models like TinyCD [7] and LightCDNet [10] that achieve superior performance while being significantly smaller and faster.
Transformer-Based CD: Transformer-based methods have emerged as a promising approach for CD. Chen et al. [23] introduced the bi-temporal image transformer (BIT), combining a transformer encoder with a ResNet backbone to model spatial-temporal contexts efficiently. Bandara et al. [12] proposed ChangeFormer, a fully transformer-based Siamese network for CD, which unifies a hierarchical transformer encoder with a multi-layer perceptron (MLP) decoder. Fang et al. [13] introduced the Changer series framework, a novel architecture for CD that incorporates alternative interaction layers between bi-temporal features. This framework is applicable to both CNN-based and Transformer-based models, enhancing the performance of the original models.
FM-Based CD: Recently, foundation models have become a new training paradigm. There have been works utilizing remote sensing data to fine-tune pretrained models such as Vision Transformers (ViT) [24], Segment Anything Model (SAM) [25], and Contrastive Language-Image Pretraining (CLIP) [26], achieving better performance in CD tasks [20, 18, 21]. Li et al. [20] proposed the Bi-Temporal Adapter Network (BAN), a universal FM-based framework for CD, which enhances existing models with minimal additional parameters and achieves significant performance improvements. Chen et al. [21] introduced Time Travelling Pixels (TTP), a method that integrates latent knowledge from the SAM model into CD, overcoming domain shifts and spatio-temporal complexities.
2.2 Knowledge Distillation in CD
3 JL1-CD Dataset
@@
Table 1: Information of open-source binary CD datasets and the proposed JL1-CD dataset.
@@
Table 1 summarizes mainstream binary CD datasets in terms of release year, number of image pairs, image size, and spatial resolution. The SZTAKI AirChange Benchmark [33] contains 12 pairs of 952 × 640 and one pair of 1,048 × 724 optical aerial images. It is one of the earliest and most widely used CD datasets in early research. The DSIFN dataset [3] consists of 6 large bi-temporal image pairs from 6 cities in China, which are cropped into 394 sub-image pairs, each sized 512 × 512. However, the resolution of all or part of the images in these datasets does not reach the sub-meter level. WHU-CD [35], LEVIR-CD [4], and S2Looking [36] are widely recognized datasets focusing on monitoring building changes. These datasets predominantly include human-induced changes and lack natural change types. The CDD dataset [34] is derived from 7 pairs of 4,725 × 2,700 real-world seasonal change remote sensing images. The SYSU-CD dataset [37] contains 20,000 pairs of 0.5-meter aerial images captured in Hong Kong between 2007 and 2014. While both datasets offer high resolution and diverse change types, the image acquisition dates are relatively outdated.
To provide a new benchmark for evaluating CD algorithms, we propose the JL1-CD dataset, a high-resolution, all-inclusive change detection dataset. JL1-CD includes 5,000 pairs of satellite images captured in China from early 2022 to the end of 2023, including Shandong, Ningxia, Anhui, Hebei, Hunan, and other regions. The images have sub-meter resolutions ranging from 0.5 to 0.75 meters and are sized 512 × 512 pixels. As shown in Figure 2, the dataset covers various common human-induced and natural surface features, such as buildings, roads, hardened surfaces, woodlands, grasslands, croplands, water bodies and photovoltaic (PV) panels. The original 5,000 image pairs are divided into 4,000 pairs for training and 1,000 pairs for testing, following a split of 80%: 20%. The JL1-CD dataset will be made openly available for all research needs.
@@
4 Methodology
4.1 O-P Strategy
@@
To quantify the diversity of changes within a dataset, we introduce the concept of Change Area Ratio (CAR), defined as the ratio of changed pixels relative to the total image area. A detailed discussion of the CAR distribution on the JL1-CD and SYSU-CD datasets is provided in Appendix A. For datasets like them, where CAR values span a wide range (0% to 100%), traditional single-stage training tends to yield models that struggle to perform well across all CAR scenarios. To address this issue, we propose a general Origin-Partition (O-P) strategy. Rather than training a single model on the entire training set, we partition the dataset into multiple subsets according to different CAR thresholds, which enables each model to focus on a specific range of change scenarios. As a concrete instantiation (see Figure  3, green boxes at Step 1 and Step 2), for a given CD algorithm, we train the corresponding models in the following sequence:
@@
As shown in the middle pink box in Figure 3, during testing, since the true CAR of the test images is unknown, we first use $\mathcal{M}_{O}$ to estimate the CAR roughly. Based on this estimated CAR, we then assign the images to one of the three categories: small, medium, or large, and send them to the corresponding model $\mathcal{M}_{T_{S}}$, $\mathcal{M}_{T_{M}}$, or $\mathcal{M}_{T_{L}}$ to obtain the final detection result:
@@
4.2 MTKD Framework
Partitioning the training set in the O-P strategy effectively reduces the training complexity for each model and can enhance CD performance to some extent. Nevertheless, this approach suffers from two major limitations. First, during inference, we are required to load four different models, and even disregarding data throughput, the time required is at least twice that of the original method, significantly increasing both memory and computational complexity. Second, since we first use $\mathcal{M}_{O}$ to obtain a rough estimate of the CAR, any inaccuracies in this estimation may lead to incorrect model selection in the subsequent steps, thereby reducing the accuracy of the final predictions. This motivates us to explore whether it is possible to integrate the strengths of the original model and those trained on sub-datasets into a single model.
To address this, we propose the Multi-Teacher Knowledge Distillation (MTKD) framework, as illustrated in the green box at the bottom of Figure 3. In the O-P strategy, we have already trained the models $\mathcal{M}_{O}$, $\mathcal{M}_{T_{S}}$, $\mathcal{M}_{T_{M}}$, and $\mathcal{M}_{T_{L}}$. Building on this, we further train a student model $\mathcal{M}_{S}$. First, we initialize the student model $\mathcal{M}_{S}$ using the parameters from $\mathcal{M}_{O}$, and then use $\mathcal{M}_{T_{S}}$, $\mathcal{M}_{T_{M}}$, and $\mathcal{M}_{T_{L}}$ as teacher models to perform KD. For each input image pair, we select the appropriate teacher model according to its CAR to supervise the student model. In this framework, the student model $\mathcal{M}_{S}$ is simultaneously supervised by the ground truth labels and the CM information from the teacher models across different CAR partitions.
@@ 4.2 MTKD Framework
4.3 Loss Function
where $Y(i)$ denotes the ground truth label of pixel $i$.
$\mathcal{L}=\mathcal{L}_{\text{CE}}+\lambda\mathcal{L}_{\text{KD}},$ (5)
Table 2: Experimental results on JL1-CD test set.
- 57.08 61.90 86.40 61.28 - 63.79 69.54 84.77 69.19
O-P 50.05 53.64 95.28 52.00 O-P 60.93 64.72 91.03 65.46
- 61.30 66.03 86.45 66.34 - 66.76 81.71 74.73 74.02
O-P 60.93 64.72 91.03 65.46 O-P 64.56 78.47 72.13 71.25
- 71.25 78.91 84.53 77.33 - 68.97 74.87 85.06 75.25
O-P 71.06 78.37 84.28 77.21 O-P 71.39 78.60 83.36 77.98
IFN MTKD 72.72 80.28 84.66 78.80 SNUNet MTKD 71.12 78.27 84.96 77.56
- 67.22 74.47 83.71 73.37 - 69.47 75.58 84.46 75.57
O-P 69.41 76.29 84.02 75.77 O-P 68.87 74.74 84.90 74.86
BIT MTKD 68.86 75.49 84.71 74.88 ChangeStar (FarSeg) MTKD 69.87 75.80 85.78 76.03
- 64.85 69.18 88.26 70.19 - 73.51 80.46 86.33 79.70
O-P 64.68 69.05 87.23 70.08 O-P 72.58 79.16 86.33 78.79
ChangeStar (UPerNet) MTKD 65.10 70.26 87.69 70.58 ChangeFormer (MiT-b0) MTKD 73.78 80.07 86.90 79.98
- 73.05 79.70 86.95 79.22 - 71.04 78.77 83.05 77.74
O-P 73.45 79.19 87.46 79.41 O-P 72.22 79.93 83.49 78.76
ChangeFormer (MiT-b1) MTKD 73.92 80.43 86.89 80.18 TinyCD MTKD 72.55 80.98 83.17 79.26
- 63.64 69.77 83.43 69.39 - 74.85 81.84 86.09 80.98
O-P 69.05 76.53 83.05 75.66 O-P 75.29 81.40 87.06 81.32
HANet MTKD 67.67 74.39 84.38 73.92 Changer (MiT-b0) MTKD 75.35 81.76 87.18 81.28
- 75.94 81.99 87.74 81.93 - 68.37 75.15 83.43 74.54
O-P 75.42 81.67 87.13 81.43 O-P 70.76 77.42 83.86 77.01
Changer (MiT-b1) MTKD 76.15 82.85 86.98 82.13 Changer (r18) MTKD 69.45 77.26 81.50 75.86
- 62.31 69.23 80.91 67.83 - 64.58 71.16 83.12 70.13
O-P 71.80 79.76 83.15 78.23 O-P 70.19 77.43 83.99 76.16
Changer (s50) MTKD 62.96 69.65 81.76 68.52 LightCDNet MTKD 65.99 72.44 83.86 71.48
- 73.37 80.31 85.33 79.65 - 73.54 79.54 87.89 79.47
O-P 72.95 79.71 85.50 79.12 O-P 73.61 79.17 88.10 79.45
CGNet MTKD 73.82 80.32 86.33 79.91 BAN MTKD 73.95 80.26 87.12 79.92
- 75.05 80.24 89.82 80.76
O-P 76.69 83.48 87.27 82.52
TTP MTKD 76.85 82.99 88.05 82.56
Figure 4: Visual comparison on the JL1-CD dataset. Each row, from left to right, represents the following: pre-event image, post-event image, ground truth, output from the original model, output from the O-P strategy, and output from the MTKD framework. Red denotes missed detections (FN), while blue indicates false alarms (FP).
Figure 5: mIoU of HANet, ChangeFormer-MiT-b1, and TTP across different CAR ranges. In each plot, the left y-axis represents CAR size, and the right y-axis represents mIoU.
5 Experiment
To comprehensively verify the validity of our JL1-CD dataset and the superiority of the proposed O-P strategy and MTKD framework, we conduct extensive experiments on existing benchmark algorithms. The benchmark methods, implementation details and evaluation metrics used in the experiments are introduced in Appendix B, C, and D , respectively.
5.1 Experimental Results
Quantitative Comparison. Table 2 summarizes the numerical results of mIoU, mRecall (mRec), mPrecision (mPrec), and mF1-score (mFscore) for all methods on the JL1-CD test set, trained under the original, O-P, and MTKD strategies. As shown in the table, the MTKD framework can improve the performance of all the algorithms to a certain extent. However, several algorithms, such as STANet, IFN, and Changer-MiT-b1, perform worse under the O-P strategy than under MTKD, and in some cases, even worse than their original models. This mainly occurs because inaccurate CAR estimation by the original model leads to the selection of an inappropriate teacher model, thereby degrading performance. More qualitative and quantitative analysis of the performance degradation caused by O-P are provided in Appendix F. This finding highlights the potential drawbacks of the O-P strategy in certain scenarios, but it also demonstrates that our MTKD method can enhance student models’ capabilities beyond those of the teacher models trained on partitioned datasets.
Visual Comparison. In remote sensing CD, four main challenges are typically encountered: false alarms, missed detections, internal density, and boundary completeness. To visually assess the effectiveness of the O-P and MTKD strategies in addressing these challenges, we present the visual results of several algorithms in Figure 4, where red indicates missed detections and blue indicates false alarms. Quantitative analysis reveals that both strategies effectively reduce false alarm and missed detection rates while improving internal consistency and boundary precision. Notably, MTKD provides greater improvements over O-P.
Results on Different CAR Partitions. We select one model from each of the CNN, Transformer, and FM architectures and evaluate their performance across different CAR ranges. Figure 5 summarizes the mIoU results of HANet, ChangeFormer-MiT-b1, and TTP on the JL1-CD test set. The images are sorted by CAR in ascending order and divided into five equal partitions. Each bar in the figure represents the lower and upper bounds of CAR for each partition, while the line graph indicates the mIoU across the different partitions. The figure indicates that the detection performance of O-P and MTKD may actually decrease for images with high CAR (e.g., ChangeFormer-MiT-b1 and TTP). However, O-P and MTKD significantly enhance model performance on images with low CAR. In the first partition of the test set (CAR range 0.02% to 0.88%), O-P improves the mIoU for the three algorithms by 8.15%, 3.21%, and 6.55%, respectively, while MTKD improves it by 4.42%, 3.36%, and 2.85%. These results show that by partitioning small CAR regions in the O-P strategy and transferring the knowledge from the corresponding teacher model to the student model in MTKD, our methods achieve substantial improvements in the accuracy of detecting tiny changes.
Comparison of Results on Change and No-Change Classes. We further select some algorithms and compare their detection performance on change and no-change regions. The experimental results are summarized in Table 4. Taking the original models as baselines, we find that MTKD yields consistently larger IoU and Fscore gains on change areas compared to no-change areas, indicating that the MTKD framework is more sensitive to the detection of change areas. For change regions, except CGNet, MTKD notably improves recall more than precision (sometimes even at the cost of reduced precision), suggesting that MTKD is particularly effective at mitigating missed detections. More experimental results on the JL1-CD dataset are provided in Appendix E.
Table 3: Comparison of detection results on change and no-change classes.
Method Class IoU Rec Prec Fscore
unchanged +0.10 -0.60 +0.65 +0.06
SNUNet changed +4.21 +7.38 -0.86 +4.54
unchanged +0.07 +0.12 -0.01 +0.05
ChangeFormer (MiT-b1) changed +1.68 +1.35 -0.11 +1.86
unchanged +0.30 +0.29 +0.08 +0.20
TinyCD changed +2.72 +4.13 +0.16 +2.85
unchanged +0.02 -0.03 -0.04 -0.06
CGNet changed +0.88 +0.03 +2.04 +0.59
unchanged +0.12 +0.23 -0.09 +0.07
BAN changed +0.70 +1.21 -1.46 +0.82
unchanged +0.23 -0.19 +0.45 +0.20
TTP changed +3.36 +5.69 -3.99 +3.39
Table 4: Impact of different numbers of teacher models.
Method Strategy $\mathcal{M}_{T}$ mIoU mFscore
3 75.29 (+0.44) 81.32 (+0.34)
O-P 2 75.44 (+0.59) 81.51 (+0.53)
3 75.35 (+0.50) 81.28 (+0.30)
Changer (MiT-b0) MTKD 2 75.72 (+0.87) 81.66 (+0.68)
3 72.95 (-0.42) 79.12 (-0.53)
O-P 2 73.56 (+0.19) 79.92 (+0.27)
3 73.82 (+0.45) 79.91 (+0.26)
CGNet MTKD 2 73.78 (+0.41) 79.89 (+0.24)
3 76.69 (+1.64) 82.52 (+1.76)
O-P 2 76.65 (+1.60) 82.49 (+1.73)
3 76.85 (+1.80) 82.56 (+1.80)
TTP MTKD 2 76.31 (+1.26) 82.22 (+1.46)
5.2 Robustness of MTKD
Impact of Different Numbers of Teacher Models on MTKD Performance. In the previous subsection, we have demonstrated that the MTKD framework based on three teacher models can effectively improve the performance of various CD models. To alleviate the resource consumption associated with training these models and testing under the O-P strategy, we further explore the performance of the MTKD framework with only two teacher models. Except for the reduced number of teacher models, all other experimental settings remain unchanged. Using the original models as baselines, Table 4 compares the metrics of models trained with the O-P strategy and the MTKD framework under different numbers of teacher models. The results indicate that two teacher models can still achieve notable performance improvements over the baseline. In certain cases, the 2-teacher configuration even outperforms the 3-teacher setup, indicating the flexibility of our approach.
Table 5: Experimental results on SYSU-CD test set.
Method Strategy mIoU mRec mPrec mFscore
- 70.13 79.20 83.41 77.80
O-P 70.25 79.28 83.88 77.81
SNUNet MTKD 70.61 79.85 83.15 78.29
- 71.19 79.29 87.43 78.41
O-P 71.76 79.84 86.62 78.80
ChangeFormer (MiT-b0) MTKD 72.16 80.63 86.63 79.44
- 76.09 84.59 87.08 82.72
O-P 75.97 84.54 86.20 82.66
TTP MTKD 76.11 83.93 87.77 82.77
Performance of MTKD on the SYSU-CD Dataset. We further evaluate the effectiveness of the MTKD framework on the SYSU-CD dataset. CNN-based SNUNet, Transformer-based ChangeFormer-MiT-b0, and FM-based TTP models are selected. As summarized in Table 5, after MTKD optimization, all the models achieve higher mIoU and mFscore values, with the most notable improvement on ChangeFormer-MiT-b0 (mIoU increased by 0.97% and mFscore increased by 1.03%). However, the O-P strategy does not yield better results on TTP compared to its original version. Overall, MTKD demonstrates substantial effectiveness on the SYSU-CD dataset, further validating its robustness. More experimental results on the SYSU-CD dataset are provided in Appendix G.
6 Conclusion
In this work, we present JL1-CD, a new benchmark dataset for CD, which substantially augments existing datasets by providing sub-meter resolution, a diverse array of change types, and a large dataset scale. The MTKD framework is further proposed, which consistently improves the performance of mainstream CD algorithms with various architectures and parameter scales, without incurring additional computational cost during inference. One potential limitation of MTKD is that the number of teacher models and the CAR thresholds are set empirically. Future work includes adaptive selection of the number of teacher models and partition thresholds based on different scenarios to further enhance the flexibility and performance of the MTKD framework.
Appendix
Appendix A CAR Distribution
Figure 6 presents sample images from the JL1-CD dataset with varying values of the Change Area Ratio (CAR). We further provide a quantitative analysis of the CAR distributions for the JL1-CD and SYSU-CD datasets in Figure 7. The CAR spans from 0 to 100% in these datasets, which poses significant challenges for model learning, and it serves as the primary motivation behind our proposed O-P strategy and MTKD framework.
(a) JL1-CD
(b) SYSU-CD
Figure 7: CAR distribution of the JL1-CD and SYSU-CD datasets.
Appendix B Benchmark Methods
To comprehensively verify the validity of our JL1-CD dataset and the effectiveness of the proposed O-P strategy and MTKD framework, we conduct extensive experiments on existing benchmark algorithms. The selected models, along with their corresponding backbones, parameter sizes, and computational complexities, are summarized in Table 6. Based on the backbone architecture, the models are categorized into three groups: Alice Blue for CNN-based models, Light Cyan for Transformer-based models, and Lavender Blue for FM-based models, encompassing almost all mainstream architectures. FLOPs is calculated with an input image of size 512 × 512. As shown in the table, the selected models span a wide range of sizes, from lightweight models with fewer than 1M parameters, such as TinyCD and LightCDNet, to the latest SOTA model TTP with over 360M parameters. This wide range of model sizes also allows us to verify the universality of the O-P and MTKD methods across models with different backbones and scales.
Table 6: Benchmark methods and the corresponding implementation details.

### v4

@@ Abstract
Change detection (CD) in remote sensing images plays a vital role in Earth observation. However, the scarcity of high-resolution, comprehensive open-source datasets and the difficulty in achieving robust performance across varying change types remain major challenges. To address these issues, we introduce JL1-CD, a large-scale, sub-meter CD dataset consisting of 5,000 image pairs. We further propose a novel Origin-Partition (O-P) strategy and integrate it into a Multi-Teacher Knowledge Distillation (MTKD) framework to enhance CD performance. The O-P strategy partitions the training set by Change Area Ratio (CAR) and trains specialized teacher models on each subset. The MTKD framework then distills complementary knowledge from these teachers into a single student model, enabling improved detection results across diverse CAR scenarios without additional inference cost. Our MTKD approach demonstrated strong performance in the 2024 “Jilin-1” Cup challenge, ranking first in the preliminary and second in the final rounds. Extensive experiments on the JL1-CD and SYSU-CD datasets show that the MTKD framework consistently improves the performance of CD models with various network architectures and parameter sizes, establishing new state-of-the-art results. Code and dataset are available at https://github.com/circleLZY/MTKD-CD.
I Introduction
@@
Figure 1: Timeline of the development of mainstream DL-based CD methods.
@@
  2. 2)
    We evaluate numerous existing algorithms on the JL1-CD dataset, establishing a new benchmark for remote sensing CD by publicly releasing the associated dataset, code and models. This contribution is anticipated to facilitate further advancements in the field.
  3. 3)
@@
II Related Works
II-A DL-Based CD
Deep learning has experienced rapid advancements, achieving remarkable success in remote sensing image CD. As illustrated in Fig. 1, we present a timeline of the development of mainstream DL-based CD algorithms. Based on the differences in neural network architectures and training paradigms, we briefly review three major categories of DL-based CD methods that are widely adopted in current research.
CNN-Based CD
CNNs serve as the foundation of many early DL-based CD methods and remain widely used today. Daudt et al. [2] proposed three fully convolutional neural network architectures, marking the entry of DL into CD research. This breakthrough spurred numerous subsequent CNN-based solutions. More recently, researchers have also developed efficient CD models like TinyCD [7] and LightCDNet [10] that achieve superior performance while being significantly smaller and faster.
Transformer-Based CD
Transformer-based methods have emerged as a promising approach for CD. Chen et al. [23] introduced the bi-temporal image transformer (BIT), combining a transformer encoder with a ResNet backbone to model spatial-temporal contexts efficiently. Bandara et al. [12] proposed ChangeFormer, a fully transformer-based Siamese network for CD, which unifies a hierarchical transformer encoder with a multi-layer perceptron (MLP) decoder. Fang et al. [13] introduced the Changer series framework, a novel architecture for CD that incorporates alternative interaction layers between bi-temporal features. This framework is applicable to both CNN-based and Transformer-based models, enhancing the performance of the original models.
FM-Based CD
Recently, foundation models have become a new training paradigm. There have been works utilizing remote sensing data to fine-tune pretrained models such as Vision Transformers (ViT) [24], Segment Anything Model (SAM) [25], and Contrastive Language-Image Pretraining (CLIP) [26], achieving better performance in CD tasks [20, 18, 21]. Li et al. [20] proposed the Bi-Temporal Adapter Network (BAN), a universal FM-based framework for CD, which enhances existing models with minimal additional parameters and achieves significant performance improvements. Chen et al. [21] introduced Time Travelling Pixels (TTP), a method that integrates latent knowledge from the SAM model into CD, overcoming domain shifts and spatio-temporal complexities.
@@ II-A DL-Based CD
TABLE I: Information of Open-Source Binary CD Datasets and the Proposed JL1-CD Dataset
@@
II-B Knowledge Distillation in CD
III JL1-CD Dataset
Table I summarizes mainstream binary CD datasets in terms of release year, number of image pairs, image size, and spatial resolution. The SZTAKI AirChange Benchmark [27] contains 12 pairs of 952 × 640 and one pair of 1,048 × 724 optical aerial images. It is one of the earliest and most widely used CD datasets in early research. The DSIFN dataset [3] consists of 6 large bi-temporal image pairs from 6 cities in China, which are cropped into 394 sub-image pairs, each sized 512 × 512. However, the resolution of all or part of the images in these datasets does not reach the sub-meter level. WHU-CD [29], LEVIR-CD [4], and S2Looking [30] are widely recognized datasets focusing on monitoring building changes. These datasets predominantly include human-induced changes and lack natural change types. The CDD dataset [28] is derived from 7 pairs of 4,725 × 2,700 real-world seasonal change remote sensing images. The SYSU-CD dataset [31] contains 20,000 pairs of 0.5-meter aerial images captured in Hong Kong between 2007 and 2014 with a size of 256 × 256. While both datasets offer high resolution and diverse change types, the image acquisition dates back several years, thus not reflecting the most recent surface changes.
To provide a new benchmark for evaluating CD algorithms, we propose the JL1-CD dataset, a high-resolution, all-inclusive change detection dataset. JL1-CD includes 5,000 pairs of satellite images captured in China from early 2022 to the end of 2023, including Shandong, Ningxia, Anhui, Hebei, Hunan, and other regions. The images have sub-meter resolutions ranging from 0.5 to 0.75 meters and are sized 512 × 512 pixels. As shown in Fig. 2, the dataset covers various common human-induced and natural surface features, such as buildings, roads, hardened surfaces, woodlands, grasslands, croplands, water bodies and photovoltaic (PV) panels. The original 5,000 image pairs are divided into 4,000 pairs for training and 1,000 pairs for testing, following a split of 80%: 20%. The JL1-CD dataset has been made openly available for all research needs.
@@ III JL1-CD Dataset
IV Methodology
In this section, we provide a comprehensive overview of the proposed methods. In Section IV-A, we first introduce the Origin-Partition (O-P) strategy designed for the challenging all-inclusive CD dataset. Building upon the O-P strategy, we further present our Multi-Teacher Knowledge Distillation (MTKD) framework in Section IV-B. Finally, in Section IV-C, we describe the overall loss function used for training the teacher and student models.
IV-A O-P Strategy
@@
To quantify the diversity of changes within a dataset, we introduce the concept of Change Area Ratio (CAR), defined as the ratio of changed pixels relative to the total image area. A detailed discussion of the CAR distribution on the JL1-CD and SYSU-CD datasets is provided in Appendix -A. For datasets like them, where CAR values span a wide range (0% to 100%), traditional single-stage training tends to yield models that struggle to perform well across all CAR scenarios. To address this issue, we propose a general Origin-Partition (O-P) strategy. Rather than training a single model on the entire training set, we partition the dataset into multiple subsets according to different CAR thresholds, which enables each model to focus on a specific range of change scenarios. As a concrete instantiation (see Fig. 3, green boxes at Step 1 and Step 2), for a given CD algorithm, we train the corresponding models in the following sequence:
@@
As shown in the middle pink box in Fig. 3, during testing, since the true CAR of the test images is unknown, we first use $\mathcal{M}_{O}$ to estimate the CAR roughly. Based on this estimated CAR, we then assign the images to one of the three categories: small, medium, or large, and send them to the corresponding model $\mathcal{M}_{T_{S}}$, $\mathcal{M}_{T_{M}}$, or $\mathcal{M}_{T_{L}}$ to obtain the final detection result:
@@
IV-B MTKD Framework
Partitioning the training set in the O-P strategy effectively reduces the training complexity for each model and can enhance CD performance to some extent. Nevertheless, this approach suffers from two major limitations:
  1. 1)
    During inference, we are required to load four different models, and even disregarding data throughput, the time required is at least twice that of the original method, significantly increasing both memory and computational complexity.
  2. 2)
    Since we first use $\mathcal{M}_{O}$ to obtain a rough estimate of the CAR, any inaccuracies in this estimation may lead to incorrect model selection in the subsequent steps, thereby reducing the accuracy of the final predictions.
These motivate us to explore whether it is possible to integrate the strengths of the original model and those trained on sub-datasets into a single model.
To address these limitations, we propose the Multi-Teacher Knowledge Distillation (MTKD) framework, as illustrated in the green box at the bottom of Fig. 3. In the O-P strategy, we have already trained the models $\mathcal{M}_{O}$, $\mathcal{M}_{T_{S}}$, $\mathcal{M}_{T_{M}}$, and $\mathcal{M}_{T_{L}}$. Building on this, we further train a student model $\mathcal{M}_{S}$. First, we initialize the student model $\mathcal{M}_{S}$ using the parameters from $\mathcal{M}_{O}$, and then use $\mathcal{M}_{T_{S}}$, $\mathcal{M}_{T_{M}}$, and $\mathcal{M}_{T_{L}}$ as teacher models to perform KD. For each input image pair, we select the appropriate teacher model according to its CAR to supervise the student model. In this framework, the student model $\mathcal{M}_{S}$ is simultaneously supervised by the ground truth labels and the CM information from the teacher models across different CAR partitions.
@@ IV-B MTKD Framework
TABLE II: Benchmark Methods and the Corresponding Implementation Details